OpenStudy (anonymous):

x+4/x-1 divided by x^2+x/x-1

OpenStudy (anonymous):

Can someone please help with this

OpenStudy (anonymous):

[(x+4)/(x-1)]/(x²+x)/(x-1) (x+4)(x-1)=(x-1)(x²+x) x²-x+4x+4=x³+x²-x²-x thus we can cancel the x² with -x², because it is null. Now we have: x²+3x+4-x³+x=0 -x³+x²+4x+4=0 to stay more simple multiplies for -1. thus: x³-x²-4x-4=0 Now is just resolve the simplified equation. I hope than helped you!!!!!!!!!!!

OpenStudy (anonymous):

(x + 4) (x - 1) --------------------- (x - 1) (x) (x + 1) x + 4 ------------- x (x + 1)